

********************************************************************
UARS WEEKLY STATUS REPORT - 19 Oct 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE resumed full operations today (Oct 19) after performing only
daylight operations since Oct 17 in support of the yaw maneuver. Otherwise,
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM was turned off in the first hour of 10/16/01; Lynn Herring is
coodinating the power on as I write on 10/17. Earlier, while
operating, SUSIM completed daily mid resolution and MgII solar scans
during periods of predicted UARS telemetry. Telemetry time, reduced
further after the official turn off, on some days is only minimally
sufficient to gather the daily mid resolution scans that are used to
make the level 3 product.
The solar pointing, recently, has been excellent. OK, folks, let's
give a nice round of applause for the project engineers, wherever they
are right now. Thank you very much.
Through devoted efforts of Charley Jackman and others, the collective
wisdom of NASA has allowed to continue to collect data in order to
test the PACOR-A level 0 SUSIM science files. Many thanks to all who
have made this possible.
***********
ACRIM
***********
The ACRIM instrument continues to operate and gather data. The last data
available from the CDHF2 was that of 9/30/01 (Mission day 3672). Data for
subsequent days is expected from the PACOR-A site but are not yet available
to us. The ACRIM2 results have been processed up through 9/30/01 and the
database is available on the acrim.com web site.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 October 2001 (Orbit 55051, GMT Day 279) through 12
October 2001 (Orbit 55156, GMT Day 285).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 279/0344 55053 HRDI Power ON
The beta angle ranged between the angles of 44.2 degrees to 23.4 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 0 degrees on 17 October 2001 (DOY 290).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.3 degrees C. Additional battery performance data for each
battery on 3 October 2001 and 15 October 2001 are provided in Appendix D.
The clock error ranged from Ð10.4 to 11.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 26
¥ HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ None this report period.
Zero (0) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 01/267 - 2287.496874
· Transponder B: 01/263 - 2287.498962
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/277 show the observatory data loss to be 303 hours, 18
minutes, and 28.016 seconds. This shows no change in data loss since the last
report period. This is a 0.39 percent data loss, which equals a 99.61 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing Raytheon's Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
MEETINGS
The UARS FOT attended a UARS Operational Plan Meeting on Wednesday, October 10,
2001.
The UARS FOT attended a UARS Network Discussion Meeting on Thursday, October
11, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 October 2001 (DOY 290). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 October 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
55051 279/0025 01-613 OBC Clk Adj fr 0.00 to 20.6 msec/day
55053 279/0344 01-614 HRDI Power ON
55054 279/0524 01-612 NBTR NOOPS D-279
55062 279/1723 01-612 NBTR NOOPS D-279
55062 279/1806 01-615 NBTR NOOPS D-280
55069 280/0427 01-615 NBTR NOOPS D-280
55072 280/0931 01-616 OBC Clk Adj fr 20.6 to 0.00 msec/day
55077 280/1724 01-615 NBTR NOOPS D-280
55077 280/1806 01-617 NBTR NOOPS D-281
55084 281/0524 01-617 NBTR NOOPS D-281
55092 281/1806 01-617 NBTR NOOPS D-282
55092 281/1806 01-618 NBTR NOOPS D-282
55099 282/0525 01-618 NBTR NOOPS D-282
55107 282/1806 01-618 NBTR NOOPS D-282
55107 282/1806 01-619 AP Update GRI
55107 282/1806 01-620 NBTR NOOPS D-283
55114 283/0525 01-620 NBTR NOOPS D-283
55122 283/1806 01-620 NBTR NOOPS D-283
55123 283/1946 01-620 NBTR NOOPS D-284
55129 284/0525 01-621 NBTR NOOPS D-284
55137 284/1806 01-621 NBTR NOOPS D-284
55137 284/1806 01-622 NBTR NOOPS D-285
55144 285/0525 01-622 NBTR NBTR PBK & RTS-58
55146 285/0714 01-622 NBTR/OBC NOOPS D-285
55152 285/1806 01-622 NBTR NOOPS D-285
55152 285/1806 01-623 NBTR NOOPS D-286
55153 285/1947 01-623 NBTR NOOPS D-286
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 October 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 October 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
55051 279/0028 Clk Adj fr 0.00 to 20.6 msec/day
55072 280/0934 Clk Adj fr 20.6 to 0.00 msec/day
__________________________________________________________________________
APPENDIX D
, UARS BATTERY PERFORMANCE DATA
Week ending 12 October 2001
03 October 2001 - GMT Day 276 Beta = 50.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/-50.0 11.7 0.000 0.0
2 +00.0/ -28.0 +4.51/+3.23 +25.2/-50.0 26.1 1.040 19.5
3 +22.4/ -22.4 +3.55/+2.29 +24.0/-50.0 26.1 1.020 19.2
15 October 2001 - GMT Day 288 Beta = 6.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/ -0.0 11.7 0.000 0.0
2 +00.0/ -95.2 +5.00/+4.19 +28.5/-12.8 25.9 1.040 19.9
3 +22.4/ -22.4 +3.87/+2.92 +27.4/-12.2 25.8 1.020 19.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 October 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
SEVNTFW01/41